# AWRaCLe: All-Weather Image Restoration using Visual In-Context Learning

### v1

@@ 1 Introduction
Unfavorable weather conditions, such as rain, snow and haze, significantly degrade the performance of computer vision systems impacting applications such as autonomous navigation, surveillance, and aerial imaging. Thus, there is a need for frameworks that mitigate weather-induced corruptions while preserving the underlying image semantics. Initial attempts at this problem (He, Sun, and Tang 2009; Roth and Black 2005; Kang, Lin, and Fu 2012) worked by modelling the physics of the degradations, and estimated physics-based parameters to obtain the restored image. These methods were not effective due to the varying degradation characteristics that occur in real-life scenarios. To overcome these limitations, many deep learning-based methods such as (Zhang, Sindagi, and Patel 2020; Zhang and Patel 2018; Wu et al. 2021; Song et al. 2023) for dehazing, (Zhang et al. 2021; Chen et al. 2020; Liu et al. 2018) for desnowing, and (Wang et al. 2019; Wei et al. 2019; Quan et al. 2021) for deraining have been proposed. However, these methods are specific to a single degradation. For handling multiple degradations, Uformer (Wang et al. 2022), Restormer (Zamir et al. 2022), MPRNet (Zamir et al. 2021) and SwinIR (Liang et al. 2021) were proposed. These approaches can be trained to remove any one type of degradation at a time. Using them for a different degradation involves either re-training or fine-tuning the networks, rendering them impractical to use in real-life settings.
@@
To address the above issues, there has been focus on developing All-Weather Image Restoration (AWIR) networks in which a single model is trained to restore multiple degradations. Popular works for tackling this problem include All-in-one (Li, Tan, and Cheong 2020), Transweather (Valanarasu, Yasarla, and Patel 2022), AirNet (Li et al. 2022), FAIG (Park, Lee, and Chun 2023), WeatherDiff (Özdenizci and Legenstein 2023), TSMC (Chen et al. 2022b), PromptIR (Potlapalli et al. 2024) and DiffUIR (Zheng et al. 2024). These methods attempt to learn degradation-specific representations from individual images during the training phase. However, they lack guidance that can provide additional information about the various degradations present in images. This limits their ability to effectively learn features unique to different degradations, ultimately impeding their performance for AWIR. Thus, there is a need for a framework capable of learning robust degradation-specific features that can facilitate effective restoration. This is quite a challenging task and proves to be difficult without supplementary knowledge about the nature of the corruption. A couple of works (Yan et al. 2023; Bai et al. 2023) have explored text-based guidance for all-weather restoration. However, text descriptions can only convey high-level semantic information about the degradation and fail to describe important aspects of the corruption such as its visual characteristics. We conjecture that feeding images as degradation-context to an image restoration network can overcome some of the above limitations as it will enable the network to aggregate visual information about different types of weather conditions. We accomplish this with the help of visual in-context learning.
@@
In-context learning, as demonstrated by large language models, is very well-studied in Natural Language Processing (NLP). In comparison, visual in-context learning is an emerging area. Providing visual context has been explored by (Bar et al. 2022), Painter (Wang et al. 2023a) and SegGPT (Wang et al. 2023b). They propose elegant solutions that involve unifying the output space of a network to solve multiple computer vision tasks based on visual context. The core idea of these frameworks is the usage of masked image modelling to extract context information from the unmasked regions of an image (as detailed in Painter) and subsequently employing in-painting for the desired prediction. Although Painter was trained for image restoration tasks using visual in-context learning, we show that it fails to use contextual information effectively for restoration tasks, thereby significantly hindering its performance. We believe this is due to two primary reasons. Firstly, the extraction of context information relies solely on the masked image modelling framework, which lacks constraints to ensure retrieval of degradation characteristics from the context images. Secondly, the context information is provided only at the input of the encoder network in the form of unmasked tokens. Consequently, there is a risk of the context information being suppressed after the initial layers of the network and even becoming negligible at the decoder.
@@
We propose AWRaCLe, a methodology for all-weather (rain, snow and haze) image restoration which elegantly leverages visual in-context learning. Our method aims to restore a query image by utilizing additional context (that we call context pair), which comprises of a degraded image and its corresponding clean version (see Fig. 1). The context requires paired images since without scene consistency, it is challenging to extract degradation-specific information (DSI) which will in-turn hamper restoration performance. In order to facilitate the restoration process, the degradation type in the context pair should align with the type of degradation present in the query image. During test time, the context pair for a degradation can be chosen from the respective training set, thus requiring only the knowledge of the type of degradation.
@@ 2.1 Adverse weather restoration
Several methods have been proposed for single weather restoration such as  (Wang et al. 2019; Wei et al. 2019) for deraining, (Zhang, Sindagi, and Patel 2020; Zhang and Patel 2018) for dehazing and  (Zhang et al. 2021; Chen et al. 2020) for desnowing. Recently, methods such as Restormer (Zamir et al. 2022) and MPRNet (Zamir et al. 2021) have tackled multiple degradations. However, the above methods require retraining or fine-tuning for each degradation. To overcome this limitation, AWIR methods have been actively explored. All-in-one (Li, Tan, and Cheong 2020) used neural architecture search to find the best-suited encoder for each degradation from a set of encoders. Transweather (Valanarasu, Yasarla, and Patel 2022) employed a unified network with a single encoder for multi-weather restoration. Airnet (Li et al. 2022) used Momentum Contrast (MoCo) (He et al. 2020) for improved degradation representations while TSMC (Chen et al. 2022b) proposed two-stage knowledge learning with multi-contrastive regularization for a similar objective. Recently, WeatherDiff (Özdenizci and Legenstein 2023) proposed a patch-based denoising diffusion model for adverse weather removal and WGWS (Zhu et al. 2023) extracted weather-general and weather-specific features for restoration. PromptIR (Potlapalli et al. 2024) utilized learnable prompt embeddings for AWIR while DiffUIR (Zheng et al. 2024) proposed selective hourglass mapping. Although several AWIR methods have been proposed, they do not utilize contextual guidance which limits their performance.
@@ 2.2 In-context Learning
Recent advancements in in-context learning have significantly improved the zero-shot performance of large language models. GPT-3 (Brown et al. 2020) demonstrated this using text completion with prompts as context while Flamingo (Alayrac et al. 2022) used language guidance for various image and video tasks. Visual in-context learning is an emerging area which is gaining increasing attention. VPI (Bar et al. 2022) proposed an image-based continuous output space for visual in-context segmentation. Directly using visual context for tackling multiple computer vision tasks is challenging due to the non-unified output space. To address this, Painter (Wang et al. 2023a) extended VPI for diverse computer vision tasks by unifying their output spaces. However, these approaches rely heavily on masked-image modeling to learn context (see Painter), which is ineffective for image restoration because there is no dedicated module to capture degradation-specific information. Additionally, context information provided at the network’s input may not propagate to deeper layers.
@@ 3.1 Degradation Context Extraction
To visualize the extracted degradation-specifc information, we overlay the output of the DCE block ($O_{\texttt{DCE}}^{l}$) for a clean-hazy and a clean-snowy context pair, respectively. This is illustrated in Fig. 3(a) where activations are overlayed on the degraded image ($I_{\texttt{d}}$) and its corresponding clean image ($I_{\texttt{c}}$). $A_{\texttt{d}}\in\mathbb{R}^{L\times C^{l}}$ and $A_{\texttt{c}}\in\mathbb{R}^{L\times C^{l}}$ represent the DCE block activations obtained by splitting $O_{\texttt{DCE}}^{l}$ for $I_{\texttt{d}}$ and $I_{\texttt{c}}$, respectively. The figure shows that the DCE block captures DSI such as the spatially-varying characteristics of haze and sparseness of snow. Furthermore, to discern this information, the DCE block uses the clean image ($I_{\texttt{c}}$) to identify and focus on degraded regions in $I_{\texttt{d}}$, evident from attention at similar locations in both $I_{\texttt{d}}$ and $I_{\texttt{c}}$. Additionally, Fig. 3(b) provides a t-SNE plot of the DCE block output ($O_{\texttt{DCE}}^{l}$), showing that DSI is clustered closely for the same type of degradation but is separated for different degradations.
@@
(b) t-SNE plot of DCE block outputs ($O_{\texttt{DCE}}^{l}$) for hazy, rainy and snowy context pairs.
@@ 3.2 Context Fusion
$I_{\texttt{q}}$ Before CF ($F^{l}_{\texttt{Proj}}$) After CF ($O_{\texttt{CF}}^{l}$)
Refer to caption Refer to caption Refer to caption
@@
Table 1: Quantitative comparisons of AWRaCLe with SOTA on the test sets described in Sec. 4.2. The values indicated are placeholders for PSNR/SSIM. Degradation type (S-Snow, R-Rain, H-Haze) is indicated within brackets. The best result is in bold, and second best is underlined.
Datasets Methods
WeatherDiff
(TPAMI’23)
WGWS
(CVPR’23)
TSMC
(CVPR’22)
AirNet
(CVPR’22)
PromptIR
(NeurIPS’23)
Painter
(CVPR’23)
DiffUIR
(CVPR’24)
AWRaCLe
SOTS Outdoor (H) 28.00/0.966 30.49/0.976 27.88/0.920 27.6/0.963 30.46/0.977 28.00/0.945 30.97/0.977 31.65/0.981
Rain100H (R) 25.84/0.824 13.91/0.410 26.48/0.822 22.96/0.692 26.32/0.821 22.50/0.792 26.49/0.788 27.20/0.840
Rain100L (R) 27.39/0.895 27.2/0.86 29.88/0.920 23.97/0.805 28.87/0.888 23.19/0.900 31.83/0.932 35.71/0.966
Snow100k (S) 31.33/0.910 32.58/0.921 32.28/0.916 29.22/0.884 33.39/0.932 27.87/0.871 31.77/0.915 33.48/0.934
Average 28.14/0.898 26.04/0.791 29.13/0.894 25.93/0.836 29.76/0.904 26.36/0.877 30.26/0.903 32.01/0.930
@@ 4.3 Comparisons
We evaluate and compare the performance of AWRaCLe with seven recent AWIR approaches on the test sets described in Sec. 4.2. The methods we use for comparison are WeatherDiff (Özdenizci and Legenstein 2023), WGWS (Zhu et al. 2023), TSMC (Chen et al. 2022b), AirNet (Li et al. 2022), PromptIR (Potlapalli et al. 2024) and DiffUIR (Zheng et al. 2024). Additionally, we also compare with Painter (Wang et al. 2023a) and show that AWRaCLe uses context much more effectively. DiffUIR is a very recent SOTA method published in CVPR 2024. For a fair comparison, WeatherDiff, WGWS, TSMC, AirNet, PromptIR, DiffUIR and Painter are retrained on the training sets mentioned in Sec. 4.2. Some recent approaches such as (Patil et al. 2023; Zhang et al. 2024; Xu et al. 2024; Ai et al. 2024) have no training code, and each method is trained on different datasets. Although the authors of (Lin et al. 2024) have released training scripts, they have not provided adequate instructions for their use. Hence, we are unable to compare with these methods. We also do not compare with methods for single weather removal as our method is proposed specifically to deal with multiple degradations.
Quantitative and Qualitative results. We discuss the performance of all the methods on the test sets described in Sec. 4.2. Table 1 contains the Peak Signal to Noise Ratio (PSNR) and Structural Similarity (SSIM) values for each method on these test sets. To evaluate AWRaCLe and Painter, we choose the context pair for each test set randomly from their respective training sets, i.e., for every test image, the context pair is chosen randomly. We resort to random selection for fairness. Moreover, the context pair is chosen from the training set, thus, requiring only the knowledge of the type of degradation during inference. From Table 1, we observe that AWRaCLe achieves excellent overall metrics. Our approach yields highest PSNR and SSIM values across all datasets. Importantly, AWRaCLe offers consistently high performance across all restoration tasks whereas competing methods perform well for some tasks but poorly for others. We also significantly outperform the in-context learning approach, Painter, highlighting the effectiveness of our in-context learning strategy. In Fig. 5, we show qualitative results for visual inspection and compare with the top-performing approaches TSMC, PromptIR and DiffUIR. It can be observed that AWRaCLe is able to handle the corruptions more effectively than the others. Additional qualitative results can be found in the supplementary along with results of AWRaCLe on real images and a discussion of the limitations of our method.
@@ 5.1 Effect of Context Pairs
Table 2: Effect of fixing a specific context pair for the test sets for different degradations.
SOTS Rain100L Rain100H Snow100k
31.61 $\pm$ 0.18 35.71 $\pm$ 0.004 27.20 $\pm$ 0.02 33.48 $\pm$ 0.01
0.981 $\pm$ $2e^{-4}$ 0.966 $\pm$ $1e^{-4}$ 0.840 $\pm$ $3e^{-4}$ 0.934 $\pm$ $3e^{-4}$

### v2

@@ 1 Introduction
Unfavorable weather conditions, such as rain, snow and haze, significantly degrade the performance of computer vision systems impacting applications such as autonomous navigation, surveillance, and aerial imaging. Thus, there is a need for frameworks that mitigate weather-induced corruptions while preserving the underlying image semantics. Intial physics-based methods (He, Sun, and Tang 2009; Roth and Black 2005; Kang, Lin, and Fu 2012) struggled to handle real-world variability in degradations. Deep learning approaches that handle a single degradation (Song et al. 2023; Chen et al. 2020; Wei et al. 2019; Liang et al. 2021; Zamir et al. 2022, 2021; Wang et al. 2022) at a time were then proposed, but they must be retrained or fine-tuned for each new condition, reducing their practicality.
@@
To address these issues, recent work has focused on All-Weather Image Restoration (AWIR) networks that handle multiple degradations with a single model (Li, Tan, and Cheong 2020; Valanarasu, Yasarla, and Patel 2022; Li et al. 2022; Park, Lee, and Chun 2023; Özdenizci and Legenstein 2023; Chen et al. 2022b; Potlapalli et al. 2024; Zheng et al. 2024). However, these approaches learn degradation representations only from individual images and lack guidance that provides detailed degradation-specific information (DSI). This limits their ability to effectively learn features unique to different degradations, impeding their performance. Thus, there is a need for a framework capable of learning robust degradation-specific features that can facilitate effective restoration. This is challenging without supplementary knowledge about the nature of the corruption. While some methods (Yan et al. 2023; Bai et al. 2023) have introduced text-based guidance for AWIR, text descriptions can only convey high-level semantic information about the degradation and fail to describe important aspects of the corruption such as its visual characteristics.
To address this limitation, recent approaches such as Diff-Plugin (Liu et al. 2024) and DA-CLIP (Luo et al. 2023) attempt to extract DSI directly from images. Diff-Plugin extracts task-specific and spatial prompts from the degraded input image. However, it requires multiple independently trained task-specific plugins for each degradation. DA-CLIP extracts text-aligned DSI from the degraded image but lacks detailed visual information, as text conveys only high-level features. Moreover, both methods face the challenge of disentangling scene characteristics from DSI because they rely on extracting DSI from a single image. We conjecture that feeding both the clean and degraded images as context to an image restoration network can overcome these limitations as the consistent scene between the pair allows the network to aggregate visual information specific to the degradation. We accomplish this with the help of visual in-context learning.
@@
In-context learning, as demonstrated by large language models, is very well-studied in Natural Language Processing (NLP). In comparison, visual in-context learning is an emerging area. Providing visual context has been explored by (Bar et al. 2022), Painter (Wang et al. 2023a), SegGPT (Wang et al. 2023b) and PromptGIP (Liu et al. 2023). They propose elegant solutions that involve unifying the output space of a network to solve multiple computer vision tasks based on visual context. The core idea of these frameworks is the usage of masked image modelling (MIM) to extract context from the unmasked regions of an image and subsequently employing in-painting for the desired prediction. Although Painter was trained for image restoration tasks using visual in-context learning, we show that it fails to use contextual information effectively, hindering its restoration performance. We believe this is due to two primary reasons. Firstly, the extraction of context relies solely on the MIM framework, which lacks constraints to ensure retrieval of degradation characteristics from the context images. Secondly, providing context only at the encoder’s input can suppress contextual information after the initial network layers, making it negligible at the decoder. PromptGIP also suffers from the above issues, leading to inferior performance.
@@
We propose AWRaCLe, a methodology for all-weather (rain, snow and haze) image restoration which elegantly leverages visual in-context learning. Our method aims to restore a query image by utilizing additional context (that we call context pair), which comprises of a degraded image and its corresponding clean version (see Fig. 1). The context requires paired images since without scene consistency, it is challenging to extract DSI which will in-turn hamper restoration performance. To facilitate restoration, the degradation type in the context pair should align with that of the query image. During test time, the context pair for a degradation can be chosen from the respective training set, thus requiring only the knowledge of the degradation type.
@@ 2.1 Adverse weather restoration
Several methods have been proposed for single weather restoration such as  (Wang et al. 2019; Wei et al. 2019) for deraining, (Zhang, Sindagi, and Patel 2020; Zhang and Patel 2018) for dehazing and  (Zhang et al. 2021a; Chen et al. 2020) for desnowing. Recently, methods such as Restormer (Zamir et al. 2022) and MPRNet (Zamir et al. 2021) have tackled multiple degradations. However, the above methods require retraining or fine-tuning for each degradation. To overcome this limitation, AWIR methods have been actively explored. All-in-one (Li, Tan, and Cheong 2020) used neural architecture search to find the best-suited encoder for each degradation from a set of encoders. Transweather (Valanarasu, Yasarla, and Patel 2022) employed a unified network with a single encoder for multi-weather restoration. Airnet (Li et al. 2022) used Momentum Contrast (MoCo) (He et al. 2020) for improved degradation representations while TSMC (Chen et al. 2022b) proposed two-stage knowledge learning with multi-contrastive regularization for a similar objective. Recently, WeatherDiff (Özdenizci and Legenstein 2023) proposed a patch-based denoising diffusion model for adverse weather removal and WGWS (Zhu et al. 2023) extracted weather-general and weather-specific features for restoration. PromptIR (Potlapalli et al. 2024) utilized learnable prompt embeddings for AWIR while DiffUIR (Zheng et al. 2024) proposed selective hourglass mapping. These AWIR methods do not utilize contextual guidance which limits their performance. Diff-Plugin (Liu et al. 2024) and DA-CLIP (Luo et al. 2023) attempt to extract degradation-specific information (DSI) directly from images to improve performance. However, this approach makes it challenging to disentangle scene characteristics from DSI, unlike our method.
@@ 2.2 In-context Learning
Recent advancements in in-context learning have significantly improved the zero-shot performance of large language models. GPT-3 (Brown et al. 2020) demonstrated this using text completion with prompts as context while Flamingo (Alayrac et al. 2022) used language guidance for various image and video tasks. Visual in-context learning is an emerging area which is gaining increasing attention. VPI (Bar et al. 2022) proposed an image-based continuous output space for visual in-context segmentation. Directly using visual context for tackling multiple computer vision tasks is challenging due to the non-unified output space. To address this, Painter (Wang et al. 2023a) extended VPI for diverse computer vision tasks by unifying their output spaces. PromptGIP (Liu et al. 2023) proposed a visual prompting question-answering framework for extracting context. However, these approaches rely heavily on masked-image modelling to learn context (see Painter), which is ineffective for image restoration because there is no dedicated module to capture degradation-specific information. Additionally, context information provided at the network’s input may not propagate to deeper layers.
@@ 3.1 Degradation Context Extraction
To visualize the extracted degradation-specifc information, we overlay the output of the DCE block ($O_{\texttt{DCE}}^{l}$) for a clean-hazy and a clean-snowy context pair, respectively. This is illustrated in Fig. 3(a) where activations are overlayed on the degraded image ($I_{\texttt{d}}$) and its corresponding clean image ($I_{\texttt{c}}$). $A_{\texttt{d}}\in\mathbb{R}^{L\times C^{l}}$ and $A_{\texttt{c}}\in\mathbb{R}^{L\times C^{l}}$ represent the DCE block activations obtained by splitting $O_{\texttt{DCE}}^{l}$ for $I_{\texttt{d}}$ and $I_{\texttt{c}}$, respectively. The figure shows that the DCE block captures DSI such as the spatially-varying characteristics of haze and sparseness of snow. Furthermore, to discern this information, the DCE block uses the clean image ($I_{\texttt{c}}$) to identify and focus on degraded regions in $I_{\texttt{d}}$, evident from attention at similar locations in both $I_{\texttt{d}}$ and $I_{\texttt{c}}$. Additionally, Fig. 3(b) provides t-SNE plots of the CLIP embeddings, $E_{\mathcal{C}}$, and DCE block outputs ($O_{\texttt{DCE}}^{l}$) for hazy, snowy and rainy context pairs. Although there is separation in the t-SNE plot with $E_{\mathcal{C}}$, it is significantly enhanced after using the DCE block. Thus, the DCE block extracts DSI that is clustered closely for the same type of degradation but is separated for different degradations.
@@
Refer to caption Refer to caption
(b) t-SNE plot of CLIP embeddings ($E_{\mathcal{C}}$, right) and DCE block outputs ($O_{\texttt{DCE}}^{l}$, left), for hazy, rainy and snowy context pairs. Separation signficantly improves after using the DCE block.
@@ 3.2 Context Fusion
$I_{\texttt{q}}$ Before CF ($F^{l}_{\texttt{Proj}}$) After CF ($O_{\texttt{CF}}^{l}$)
Refer to caption Refer to caption Refer to caption
@@
Table 1: Quantitative comparisons of AWRaCLe with SOTA on the test sets described in Sec. 4.2. The values indicated are placeholders for PSNR/SSIM. The best result is in bold, and second best is underlined.
Datasets WeatherDiff WGWS TSMC AirNet PromptIR Painter DA-CLIP DiffUIR DiffPlugin PromptGIP AWRaCLe
TPAMI’23 CVPR’23 CVPR’22 CVPR’22 NeurIPS’23 CVPR’23 ICLR’24 CVPR’24 CVPR’24 ICML’24 -
SOTS 28.0/0.966 30.5/0.976 27.9/0.920 27.6/0.963 30.5/0.977 28.0/0.945 26.9/0.958 31.0/0.977 23.6/0.778 17.9/0.672 31.7/0.981
Rain100H 25.8/0.824 13.9/0.410 26.5/0.822 23.0/0.692 26.3/0.821 22.5/0.792 23.4/0.730 26.5/0.788 16.1/0.527 18.0/0.482 27.2/0.840
Rain100L 27.4/0.895 27.2/0.860 29.9/0.920 24.0/0.805 28.9/0.888 23.2/0.900 30.1/0.918 31.8/0.932 25.4/0.698 22.8/0.662 35.7/0.966
Snow100k 31.3/0.910 32.6/0.921 32.3/0.916 29.2/0.884 33.4/0.932 27.9/0.871 30.6/0.893 31.8/0.915 23.5/0.658 20.8/0.615 33.5/0.934
Average 28.1/0.898 26.0/0.791 29.1/0.894 25.9/0.836 29.8/0.904 26.4/0.877 27.8/0.874 30.3/0.903 22.2/0.665 19.9/0.608 32.0/0.930
@@ 4.3 Comparisons
We evaluate and compare the performance of AWRaCLe with ten recent AWIR approaches on the test sets described in Sec. 4.2. The methods we use for comparison are WeatherDiff (Özdenizci and Legenstein 2023), WGWS (Zhu et al. 2023), TSMC (Chen et al. 2022b), AirNet (Li et al. 2022), PromptIR (Potlapalli et al. 2024), DiffUIR (Zheng et al. 2024), DiffPlugin (Liu et al. 2024) and DA-CLIP (Luo et al. 2023). Additionally, we also compare with Painter (Wang et al. 2023a) and PromptGIP (Liu et al. 2023), and show that AWRaCLe uses context much more effectively. For a fair comparison, all methods are retrained on the training sets mentioned in Sec. 4.2. Some recent approaches such as (Patil et al. 2023; Zhang et al. 2024; Xu et al. 2024; Ai et al. 2024) have no training code. Hence, we are unable to compare with these methods. We also do not compare with methods for single weather removal as our method is proposed specifically to deal with multiple degradations.
Quantitative and Qualitative results. We discuss the performance of all the methods on the test sets described in Sec. 4.2. Table 1 contains the Peak Signal to Noise Ratio (PSNR) and Structural Similarity (SSIM) values for each method on these test sets. To evaluate AWRaCLe, Painter and PromptGIP, we choose the context pair for each test set randomly from their respective training sets, i.e., for every test image, the context pair is chosen randomly. We resort to random selection for fairness. Moreover, the context pair is chosen from the training set, thus, requiring only the knowledge of the type of degradation during inference. From Table 1, we observe that AWRaCLe achieves excellent overall metrics. Our approach yields highest PSNR and SSIM values across all datasets. Importantly, AWRaCLe offers consistently high performance across all restoration tasks whereas competing methods perform well for some tasks but poorly for others. We also significantly outperform the in-context learning approaches, Painter and PromptGIP, highlighting the effectiveness of our in-context learning strategy. Additionally, we provide quantitative comparisons with LPIPS and FID scores in the supplementary.
In Fig. 5, we show qualitative results for visual inspection and compare with the top-performing approaches TSMC, PromptIR and DiffUIR. It can be observed that AWRaCLe is able to handle the corruptions more effectively than the others. More qualitative results, performance of AWRaCLE on real images along with a user study, and a discussion of the limitations of our method are provided in the supplementary.
@@ 5 Ablation Studies
Table 2: Effect of fixing a specific context pair for the test sets described in Sec. 4.2 for different degradations.
SOTS Rain100L Rain100H Snow100k
[…]
@@ 5.1 Effect of Context Pairs
Finally, we tested our model’s robustness to out-of-distribution (OoD) context pairs by sampling them from the following datasets unseen during training: Foggy Cityscapes (Hahner et al. 2019) for dehazing, rain images from RainDS (Quan et al. 2021) for deraining, and SnowCityscapes (Zhang et al. 2021b) for desnowing. Our model obtained PSNR/SSIM of $31.02/0.976$ on SOTS Outdoor, $35.72/0.966$ on Rain100L, $27.19/0.840$ on Rain100H and $33.40/0.934$ on Snow100k datasets. These results show only minimal deviations from those reported in Table 1, showcasing our model’s resilience to out-of-distribution context.